### v1

Abstract
Digital health tools have the potential to significantly improve the delivery of healthcare services. However, their use remains comparatively limited due, in part, to challenges surrounding usability and trust. Recently, Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare. Directly applying LLMs in clinical settings is not straightforward, with LLMs susceptible to providing inconsistent or nonsensical answers. We demonstrate how LLMs can utilize external tools to provide a novel interface between clinicians and digital technologies. This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLM in clinical settings such as hallucinations. We illustrate our approach with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
Digital healthcare technologies represent a frontier in medicine. A multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis . Clinical adoption of such methods, however, faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise . One key issue is usability Ratwani2019decade , which can result in errors associated with patient harm Howe2018electronic and contribute to clinician frustration, jeopardizing patient safety Shanafelt2016 ; gardner2019physician .
New tools employing artificial intelligence (AI) and machine learning offer substantial promise Bajwae2021 . Yet these approaches face the same usability challenges as existing digital tools, while AI systems present new questions around model trust Rajpurkar2022 ; asan2020artificial . These issues have limited the uptake and impact of AI technologies thus far goldfarb2022ai ; Davenport2019 .
Large Language Models (LLMs) have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction. Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 . While LLMs have general capabilities, they can lack domain-specific knowledge. This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 .
Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores. While others have discussed LLMs replacing existing clinical tools, we envisage LLMs forming a novel interface with clinical tools by providing a dynamic output and facilitating deeper interaction with the tools and related information, such as model explainability, medical papers, and clinical guidelines.
In addition to generating text, LLMs can also make predictions from text data. For example, Jiang et al. trained an LLM to predict hospital readmission, in-hospital mortality, and insurance denial, demonstrating the additional information in medical notes compared to structured features Jiang2023health . However, most data in medicine is not text and, consequently, other predictive models are often preferred hegselmann2023tabllm .
Applying LLMs in safety-critical clinical settings is not straightforward. LLMs may provide inconsistent or nonsensical answers Singhal2023 ; Lecler2023revolutionizing and have a tendency to hallucinate facts maynez2020faithfulness ; Ji2023survey . This is clearly unacceptable in medicine when making high-stakes decisions. Additionally, LLMs can encounter difficulty with seemingly basic functionality, such as mathematical calculations or factual lookup patel2021nlp ; schick2023toolformer , and are unable to access up-to-date information komeili2022internet . These limitations constrain the clinical utility of LLMs.
Table 1: Utility of LLMs as an interface for risk scoring. Representative questions that a clinician might have of a risk score at different stages, together with whether existing interfaces for risk scores provide this information. All questions can be addressed using LLM-based interfaces.
Existing
Stage Representative questions interfaces
Before Patient
Which features does the risk score use?
Encounter
Why are these features included in the risk score?
How was the risk score validated?
What is the methodology underlying the risk score?
Before Risk
When do clinical guidelines recommend risk scoring?
Scoring
What is the recommended risk score?
Who is the risk score suitable for?
During Risk
What is the risk for this patient?
Scoring
What characteristics led to the patient’s risk?
What effect would changing this feature have on the risk?
After Risk Scoring
What action is recommended by the guidelines based on the risk?
Despite these challenges, LLMs offer a potential solution to the challenges faced by digital tools and risk scores. Specifically, we propose using LLMs as comprehension engines for mediating human-AI interaction. We show how LLMs can form unique interfaces between healthcare professionals and clinical predictive models. By replacing existing fixed user interfaces or application programming interfaces (APIs), LLMs enable dynamic interactions that can streamline digital workflows and improve efficiency and usability (Fig. 1). For example, numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1). Existing interfaces only serve to calculate risk. Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines.
We achieve this by enabling LLMs to use external tools, thereby not solely relying on the inherent capabilities of a given LLM but imbuing them with approved medical tools and other sources of information. This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface. This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and is in contrast to previous applications of LLMs in medicine. LLMs help engender trust in risk scores by enabling increased, dynamic interactions with the model, previously identified as a key feature lakkaraju2022rethinking . The ability to adapt and tailor interactions is important since the most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient. Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment. From an LLM perspective, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
1 Illustrative LLM Interfaces
We exemplify our approach with applications to risk prediction models for cardiovascular disease (CVD) and diabetes. Our approach is general, requires no additional training, and scales to several diverse tools. To demonstrate the versatility of our approach, we take an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbue it with a set of tools using in-context learning dong2023survey . Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool and, if so, which tool and with what input. This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source.
CVD Risk Prediction. We first show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 2). We provide the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD . Additionally, we provide the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD . An example interaction between a physician and the LLM-based system is shown in Fig. 1(a). An illustration of the reasoning process by which the LLM uses external tools is provided in Fig. 1(b).
The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD. The LLM then summarizes the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines. Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming they were younger. While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 . In the following example, we will explore a more dynamic output extending beyond model use and information retrieval to new functionality.
Figure 2: (a) Example interaction with an LLM-based interface for CVD. The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and finally provides the recommended action for this patient from the NICE clinical guidelines NICE_CVD . (b) Illustration of the process by which the LLM uses external tools.
Diabetes Risk Prediction. We now demonstrate an application for diabetes risk prediction using a machine learning model (Fig. 3). We train a prognostic model using AutoPrognosis Imrie2023AutoPrognosis on a cohort from UK Biobank Sudlow2015 to predict 10-year risk of developing diabetes.
Figure 3: Example interaction with an LLM for diabetes. After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient. An example interaction is shown in Fig. 3.
After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model. Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions. In the example, the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified , but the LLM is able to explain how it works. This demonstrates the utility of LLMs beyond simply using existing tools and illustrates the benefit of their underlying knowledge.
2 Discussion
Large Language Models hold substantial promise for the medical domain, particularly in augmenting digital workflows and improving efficiency in healthcare delivery. The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as the potential to hallucinate.
We have demonstrated how LLMs can provide a unique interface between healthcare professionals and clinical predictive models, such as risk scores. Our approach is scalable and requires no additional training, although approaches that improve with use could be yet more powerful. Additionally, we aim to mitigate the problem of LLM hallucination and ensure that actionable advice is anchored in approved clinical resources.
As AI in medicine continues to advance, further research into LLMs and their potential applications is not only beneficial but vital. LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes. While we believe this paper represents an important first step, we are only scratching the surface of the potential of LLMs in healthcare. Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
3 Methods
3.1 LLMs
We used GPT-4 openai2023gpt4 as a pretrained LLM, with access through the OpenAI API. To enable the model to use external tools, we use the ReAct framework yao2022react implemented using LangChain LangChain . Toolformer taught LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning schick2023toolformer . We used in-context learning dong2023survey to remove the need for further training of the LLM, which might be challenging in the medical domain without suitable examples. This also enabled us to use multiple tools. We built user interfaces using StreamLit streamlit .
3.2 Prognostic modeling
We used QRisk3 Hippisley-Cox2017CVD to assess the likelihood of developing CVD within 10 years. For the diabetes risk score, we trained a predictive model using AutoPrognosis 2.0 Imrie2023AutoPrognosis on a cohort of individuals from UK Biobank Sudlow2015 with no known history of diabetes at baseline. The primary outcome was diagnosis of diabetes within a 10-year horizon. We selected 20 features based on their predictive value Imrie2023AutoPrognosis , namely HbA1c, blood glucose, sex, age, waist size, hip size, weight, BMI, waist-height ratio, waist-hip ratio, gamma-glutamyl transferase, cystatin-c, C-reactive protein, atrial fibrillation, HDL, alanine transaminase, triglycerides, uric acid, sex hormone-binding globulin, use of anti-hypertensives. To interpret model predictions, we used SHAP lundberg2017unified .
4 Data availability
This research has been conducted using the UK Biobank resource. Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register). The authors had no special access or privileges when accessing the data.
[…]
Acknowledgments The authors received no specific funding for this work.
[…]
Competing interests The authors declare no competing interests.

### v2

@@
 Digital health tools have the potential to significantly improve the delivery of healthcare services.
-However, their use remains comparatively limited due, in part, to challenges surrounding usability and trust.
-Recently, Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare.
-Directly applying LLMs in clinical settings is not straightforward, with LLMs susceptible to providing inconsistent or nonsensical answers.
-We demonstrate how LLMs can utilize external tools to provide a novel interface between clinicians and digital technologies.
-This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLM in clinical settings such as hallucinations.
-We illustrate our approach with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
-Digital healthcare technologies represent a frontier in medicine.
-A multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis .
+However, their adoption remains comparatively limited due, in part, to challenges surrounding usability and trust.
+Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare.
+Directly applying LLMs in clinical settings is not straightforward, however, with LLMs susceptible to providing inconsistent or nonsensical answers.
+We describe how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies.
+This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLMs in clinical settings such as hallucinations.
+We illustrate LLM-based interfaces with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
+Digital healthcare technologies represent a frontier in medicine and a multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis .
 Clinical adoption of such methods, however, faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise .
@@
 New tools employing artificial intelligence (AI) and machine learning offer substantial promise Bajwae2021 .
-Yet these approaches face the same usability challenges as existing digital tools, while AI systems present new questions around model trust Rajpurkar2022 ; asan2020artificial .
+Yet these approaches face the same usability challenges as existing digital tools, while AI systems present additional questions around model trust Rajpurkar2022 ; asan2020artificial .
 These issues have limited the uptake and impact of AI technologies thus far goldfarb2022ai ; Davenport2019 .
-Large Language Models (LLMs) have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction.
+Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools and risk scores if used appropriately.
+LLMs have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction.
 Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 .
-While LLMs have general capabilities, they can lack domain-specific knowledge.
-This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 .
 Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores.
 While others have discussed LLMs replacing existing clinical tools, we envisage LLMs forming a novel interface with clinical tools by providing a dynamic output and facilitating deeper interaction with the tools and related information, such as model explainability, medical papers, and clinical guidelines.
-In addition to generating text, LLMs can also make predictions from text data.
-For example, Jiang et al.
-trained an LLM to predict hospital readmission, in-hospital mortality, and insurance denial, demonstrating the additional information in medical notes compared to structured features Jiang2023health .
-However, most data in medicine is not text and, consequently, other predictive models are often preferred hegselmann2023tabllm .
 Applying LLMs in safety-critical clinical settings is not straightforward.
@@
 This is clearly unacceptable in medicine when making high-stakes decisions.
-Additionally, LLMs can encounter difficulty with seemingly basic functionality, such as mathematical calculations or factual lookup patel2021nlp ; schick2023toolformer , and are unable to access up-to-date information komeili2022internet .
+Additionally, LLMs can encounter difficulty with seemingly basic functionality, such as mathematical calculations or factual lookup patel2021nlp ; schick2023toolformer , and are unable to access up-to-date information by default komeili2022internet .
 These limitations constrain the clinical utility of LLMs.
-Table 1: Utility of LLMs as an interface for risk scoring.
-Representative questions that a clinician might have of a risk score at different stages, together with whether existing interfaces for risk scores provide this information.
+In this Comment, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools.
+Conceptually, this differs substantially from previous applications, such as training medicine-specific LLMs Luo2022 ; Yang2022large or using LLMs for prediction Jiang2023health .
+Instead, we construct LLM-based systems, offering an intuitive natural language interface that can streamline clinician interactions with multiple digital tools and sources of information, improving efficiency and usability (Fig.
+1).
+By adopting a systems approach, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
+LLM-based user interfaces can help engender trust in digital tools and AI models, such as risk scores, by enabling dynamic interactions in the form of natural language dialogues, which has been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking .
+The most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient.
+Thus, the ability to adapt and tailor interactions represents a critical advance in the functionality of such tools that can be unlocked by LLMs.
+Table 1: Representative questions that a clinician might have relating to a risk score at different stages of a patient encounter, together with whether existing interfaces for risk scores provide this information.
 All questions can be addressed using LLM-based interfaces.
@@
 What action is recommended by the guidelines based on the risk?
-Despite these challenges, LLMs offer a potential solution to the challenges faced by digital tools and risk scores.
-Specifically, we propose using LLMs as comprehension engines for mediating human-AI interaction.
-We show how LLMs can form unique interfaces between healthcare professionals and clinical predictive models.
-By replacing existing fixed user interfaces or application programming interfaces (APIs), LLMs enable dynamic interactions that can streamline digital workflows and improve efficiency and usability (Fig.
-1).
-For example, numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1).
-Existing interfaces only serve to calculate risk.
-Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines.
-We achieve this by enabling LLMs to use external tools, thereby not solely relying on the inherent capabilities of a given LLM but imbuing them with approved medical tools and other sources of information.
+As a concrete example, let us consider risk scores.
+Currently, clinicians access these tools via fixed user interfaces or application programming interfaces (APIs).
+Numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1), while existing interfaces only serve to calculate risk.
+Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
+By default, LLMs do not possess the ability to access external tools or information.
+Methods to extend the functionality of LLMs beyond text generation are in their infancy but already enable powerful LLM-based systems.
+Below, we describe such a system, illustrated in Fig.
+2.
+Instead of solely relying on an LLM to issue predictions or provide information, we can augment LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM.
 This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface.
-This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and is in contrast to previous applications of LLMs in medicine.
-LLMs help engender trust in risk scores by enabling increased, dynamic interactions with the model, previously identified as a key feature lakkaraju2022rethinking .
-The ability to adapt and tailor interactions is important since the most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient.
-Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment.
-From an LLM perspective, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
+This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and contrasts previous applications of LLMs in medicine.
 1 Illustrative LLM Interfaces
-We exemplify our approach with applications to risk prediction models for cardiovascular disease (CVD) and diabetes.
-Our approach is general, requires no additional training, and scales to several diverse tools.
-To demonstrate the versatility of our approach, we take an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbue it with a set of tools using in-context learning dong2023survey .
-Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool and, if so, which tool and with what input.
-This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source.
+To illustrate the ideas described in this paper, we built example instantiations of LLM-based interfaces for cardiovascular disease (CVD) and diabetes risk prediction (Fig.
+3).
+To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM.
+Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input.
+This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM.
+Additional implementation details are provided in Appendix A.
 CVD Risk Prediction.
 We first show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig.
-2).
-We provide the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
-QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD .
-Additionally, we provide the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
+2(a)).
+We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
+Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
 An example interaction between a physician and the LLM-based system is shown in Fig.
-1(a).
+2(a).
 An illustration of the reasoning process by which the LLM uses external tools is provided in Fig.
-1(b).
-The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD.
+S.1).
+Figure 2: Overview of an LLM-based system that enables clinicians to interface with digital tools using natural language inputs.
+(1) The LLM is provided with the history of the interaction, including the current request.
+(2) Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input.
+(3) The LLM provides a response to the user incorporating information provided by any tools that were used.
+In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM uses QRisk3 to estimate the patient’s risk, providing this to the user.
 The LLM then summarizes the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines.
-Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming they were younger.
+Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges.
+This allows both the clinician and patient to understand the potential impact of changes to modifiable variables on the patient’s risk.
 While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 .
 In the following example, we will explore a more dynamic output extending beyond model use and information retrieval to new functionality.
-Figure 2: (a) Example interaction with an LLM-based interface for CVD.
-The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and finally provides the recommended action for this patient from the NICE clinical guidelines NICE_CVD .
-(b) Illustration of the process by which the LLM uses external tools.
+(a) Cardiovascular disease risk prediction
+(b) Diabetes risk prediction
+Figure 3: Example interactions with LLM-based interfaces.
+(a) CVD: The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines.
+(b) Diabetes: After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
 Diabetes Risk Prediction.
 We now demonstrate an application for diabetes risk prediction using a machine learning model (Fig.
-3).
-We train a prognostic model using AutoPrognosis Imrie2023AutoPrognosis on a cohort from UK Biobank Sudlow2015 to predict 10-year risk of developing diabetes.
-Figure 3: Example interaction with an LLM for diabetes.
-After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
-Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial .
+2(b)).
+We trained a prognostic model using AutoPrognosis Imrie2023AutoPrognosis , an open-source automated machine learning framework, on a cohort of individuals from UK Biobank Sudlow2015 to estimate a patient’s 10-year risk of developing diabetes.
+Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 .
 A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
-To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient.
-An example interaction is shown in Fig.
-3.
-After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model.
-Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions.
-In the example, the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified , but the LLM is able to explain how it works.
+To enable this, in addition to allowing the LLM access to the model, we provided the LLM with the ability to apply explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions.
+As shown in Fig.
+2(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model.
+In this example, we demonstrate the use of Shapley additive explanations (SHAP) lundberg2017unified .
+Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM.
+For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified .
+The LLM is able to explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified.
 This demonstrates the utility of LLMs beyond simply using existing tools and illustrates the benefit of their underlying knowledge.
@@
 Additionally, we aim to mitigate the problem of LLM hallucination and ensure that actionable advice is anchored in approved clinical resources.
+While LLMs have general capabilities, they can lack domain-specific knowledge.
+This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 .
+While we showed using such specialist LLMs is not required, they are easy to incorporate due to the modularity of the approach and doing so could further enhance the functionality of LLM-based interfaces.
 As AI in medicine continues to advance, further research into LLMs and their potential applications is not only beneficial but vital.
 LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes.
+Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment.
 While we believe this paper represents an important first step, we are only scratching the surface of the potential of LLMs in healthcare.
 Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
-3 Methods
-3.1 LLMs
-We used GPT-4 openai2023gpt4 as a pretrained LLM, with access through the OpenAI API.
-To enable the model to use external tools, we use the ReAct framework yao2022react implemented using LangChain LangChain .
-Toolformer taught LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning schick2023toolformer .
-We used in-context learning dong2023survey to remove the need for further training of the LLM, which might be challenging in the medical domain without suitable examples.
+3 Code availability
+Code for the LLM-based interfaces can be accessed at https://github.com/pauliusrauba/LLMs_interface.
+4 Data availability
+The risk predictor for diabetes was trained on a cohort from the UK Biobank resource.
+Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register).
+The authors had no special access or privileges when accessing the data.
+[…]
+Acknowledgments This study received no funding.
+[…]
+Competing interests All authors declare no financial or non-financial competing interests.
+[…]
+Author contributions F.I.
+and M.vdS.
+conceptualized the manuscript.
+F.I.
+and P.R.
+designed and performed the experiments.
+F.I.
+wrote the original draft of the manuscript, and all authors contributed to editing and revising it.
+    A.1 Large Language Models
+    By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history.
+To construct interfaces for digital tools using LLMs, we instead view the LLM as an agent that can interact with an environment to solve tasks.
+Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history.
+To enable the agent to both reason and use external tools, we use the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
+    Since we will provide the agents tasks in the form of natural language, and actions in the language space $L$ are (essentially) infinite, we chose to benefit from strong language priors and use a pretrained LLM.
+We implemented our LLM-based interfaces using GPT-4 openai2023gpt4 , with access through the OpenAI API.
+Interactions with external tools were implemented using LangChain LangChain .
+    While frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively.
+Following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM.
+This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples.
 This also enabled us to use multiple tools.
 We built user interfaces using StreamLit streamlit .
-3.2 Prognostic modeling
-We used QRisk3 Hippisley-Cox2017CVD to assess the likelihood of developing CVD within 10 years.
+    A.2 External Tools
+    This section describes the external tools and sources of information made available to the LLM in our two examples. In the subsequent section, we provide the system and tool prompts.
+    CVD Risk Prediction.
+In our first example, we show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig.
+3a).
+We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
+QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD .
+We enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk.
+Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
+    Diabetes Risk Prediction.
+To demonstrate an application for diabetes risk prediction using a machine learning model (Fig.
+3b), we equipped the LLM with the following tools.
 For the diabetes risk score, we trained a predictive model using AutoPrognosis 2.0 Imrie2023AutoPrognosis on a cohort of individuals from UK Biobank Sudlow2015 with no known history of diabetes at baseline.
@@
 We selected 20 features based on their predictive value Imrie2023AutoPrognosis , namely HbA1c, blood glucose, sex, age, waist size, hip size, weight, BMI, waist-height ratio, waist-hip ratio, gamma-glutamyl transferase, cystatin-c, C-reactive protein, atrial fibrillation, HDL, alanine transaminase, triglycerides, uric acid, sex hormone-binding globulin, use of anti-hypertensives.
-To interpret model predictions, we used SHAP lundberg2017unified .
-4 Data availability
-This research has been conducted using the UK Biobank resource.
-Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register).
-The authors had no special access or privileges when accessing the data.
-[…]
-Acknowledgments The authors received no specific funding for this work.
-[…]
-Competing interests The authors declare no competing interests.
+    Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial .
+A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
+To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient.
+In particular, we used SHAP lundberg2017unified to interpret model predictions.
+    Figure S.1: Illustration of the process by which the LLM uses external tools.
+The LLM is provided with the history of the interaction, including the current request.
+Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input (“Thought”).
+The LLM then uses the external tool (“Action”) and receives the output (“Observation”).
+Finally, the LLM decides to answer the question (“Thought”), providing a response to the user.

### v3

@@
 Directly applying LLMs in clinical settings is not straightforward, however, with LLMs susceptible to providing inconsistent or nonsensical answers.
-We describe how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies.
+We demonstrate how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies.
 This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLMs in clinical settings such as hallucinations.
-We illustrate LLM-based interfaces with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
-Digital healthcare technologies represent a frontier in medicine and a multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis .
-Clinical adoption of such methods, however, faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise .
+We illustrate LLM-based interfaces with the example of cardiovascular disease risk prediction.
+We develop a new prognostic tool using automated machine learning and demonstrate how LLMs can provide a unique interface to both our model and existing risk scores, highlighting the benefit compared to traditional interfaces for digital tools.
+1 Introduction
+Digital healthcare technologies represent a frontier in medicine.
+Despite a multitude of tools being developed sutton2020overview ; dunn2018wearables , clinical adoption of such methods faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise .
 One key issue is usability Ratwani2019decade , which can result in errors associated with patient harm Howe2018electronic and contribute to clinician frustration, jeopardizing patient safety Shanafelt2016 ; gardner2019physician .
-New tools employing artificial intelligence (AI) and machine learning offer substantial promise Bajwae2021 .
-Yet these approaches face the same usability challenges as existing digital tools, while AI systems present additional questions around model trust Rajpurkar2022 ; asan2020artificial .
-These issues have limited the uptake and impact of AI technologies thus far goldfarb2022ai ; Davenport2019 .
-Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools and risk scores if used appropriately.
+New tools employing artificial intelligence (AI) and machine learning offer substantial promise, with their impact expected to be felt across all areas of healthcare Bajwae2021 .
+Yet these approaches face the same usability challenges as existing digital tools, while presenting additional questions around model trust Rajpurkar2022 ; asan2020artificial .
+Consequently, these issues have limited the uptake and impact of AI technologies in clinical settings thus far goldfarb2022ai ; Davenport2019 ; Kelly2019 .
+To improve the usability of clinical predictive models, several approaches have sought to simplify or automate the process of obtaining a prediction.
+These include points-based scoring systems Gage2001 , web-based calculators Hippisley-Cox2017CVD ; Imrie2023AutoPrognosis , or integration within electronic health records Rothman2013 .
+While this can make such tools easier to use, simply obtaining a prediction is frequently insufficient and more is required to build model trust with both clinicians Rajpurkar2022 and regulators FDA2019 ; Mourby2021 .
+Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools.
 LLMs have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction.
 Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 .
-Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores.
-While others have discussed LLMs replacing existing clinical tools, we envisage LLMs forming a novel interface with clinical tools by providing a dynamic output and facilitating deeper interaction with the tools and related information, such as model explainability, medical papers, and clinical guidelines.
 Applying LLMs in safety-critical clinical settings is not straightforward.
@@
 These limitations constrain the clinical utility of LLMs.
-In this Comment, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools.
+In this paper, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools.
 Conceptually, this differs substantially from previous applications, such as training medicine-specific LLMs Luo2022 ; Yang2022large or using LLMs for prediction Jiang2023health .
@@
 1).
-By adopting a systems approach, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
-LLM-based user interfaces can help engender trust in digital tools and AI models, such as risk scores, by enabling dynamic interactions in the form of natural language dialogues, which has been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking .
+Dynamic interactions in the form of natural language dialogues have been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking .
 The most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient.
 Thus, the ability to adapt and tailor interactions represents a critical advance in the functionality of such tools that can be unlocked by LLMs.
+By default, LLMs do not possess the ability to access external tools or information.
+We augment the base functionality of LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM.
+This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface.
+By adopting a systems approach, the LLM does not itself issue predictions and can access relevant domain-specific information, rather than needing to supply specific knowledge itself.
+Consequently, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
+Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores.
+While others have discussed LLMs replacing existing clinical tools (1,2), we envisage LLMs forming a novel interface by enabling dynamic interactions and facilitating deeper engagement with tools and related information, such as model explainability, medical papers, and clinical guidelines (3).
+As an example of our approach, we specifically focused on risk scoring moons2009prognosis and considered cardiovascular disease (CVD), the most common cause of mortality globally Muthiah2022 .
+Primary prevention uses prognostic models to estimate the future risk of developing CVD.
+This allows high-risk individuals to be identified and their risk to be managed via interventions, such as lifestyle modifications or pharmaceuticals.
+Numerous CVD risk scores have previously been developed, for example the Framingham score DAgostino2008 in the United States, SCORE2 SCORE2 in Europe, and QRisk3 Hippisley-Cox2017CVD in the UK.
+We begin by developing an improved risk score for CVD using AutoPrognosis 2.0 Imrie2023AutoPrognosis , an open-source automated machine learning framework for developing diagnostic and prognostic models.
+We then demonstrate a novel LLM-based interface for both our prognostic model and existing risk scores.
+We provide several illustrative examples of dynamic interactions that substantially extend the capabilities of existing fixed interfaces.
+Table 1: Model validation on UK Biobank.
+C-indices, Brier scores, and Expected/Observed ratios for our approach (AP2), QRisk3, SCORE2, and Framingham score.
+We report results for the original existing risk scores and the performance following recalibration.
+Mean performance (95% CI).
+C-Index $\uparrow$ Brier Score $\downarrow$ E/O
+AP2 0.741 (0.740-0.743) 0.041 (0.041-0.041) 1.003 (1.001-1.006)
+Original Risk Scores
+Framingham 0.705 (0.703-0.707) 0.047 (0.047-0.047) 2.495 (2.491-2.499)
+SCORE2 0.710 (0.709-0.712) 0.042 (0.042-0.042) 1.606 (1.603-1.608)
+QRisk3 0.722 (0.721-0.724) 0.044 (0.043-0.044) 1.941 (1.937-1.944)
+Recalibrated Risk Scores
+Framingham 0.705 (0.703-0.707) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
+SCORE2 0.710 (0.709-0.712) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
+QRisk3 0.722 (0.721-0.724) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
+2 Results
+We conducted experiments using data from the UK Biobank Sudlow2015 , a prospective population study of around half a million individuals from the UK, enrolled between 2006 and 2010.
+UK Biobank collected a broad set of information from participants at enrollment, including health and medical history, blood tests, physical examination, and socio-demographics, with ongoing linkage to healthcare records.
+The outcome of interest was the incidence of CVD within a 10-year horizon, where CVD was defined as the composite endpoint of myocardial infarction (ICD-10 codes: I21, I22), angina pectoris (I20), stroke (I63, I64), or transient cerebral ischaemic attacks (G45).
+The cohort contained 407,605 individuals and there were 17,600 CVD events within a 10-year horizon.
+A flow chart of the study cohort is provided in Fig.
+S.2 and a summary of patient characteristics is provided in Table S.1.
+Further details can be found in Section 4.
+We used AutoPrognosis 2.0 Imrie2023AutoPrognosis to automatically construct a prognostic model for incidence of CVD.
+We first validate the performance of our approach compared to existing risk scores and then demonstrate a novel interface for using such risk scores using LLMs.
+2.1 Model performance
+In Table 1, we compared our prognostic model (AP2) with three existing CVD risk scores, namely QRisk3 Hippisley-Cox2017CVD , SCORE2 SCORE2 , and Framingham score DAgostino2008 .
+The ensemble model constructed using AutoPrognosis achieved a C-index of 0.741 (95% CI: 0.740-0.743), a Brier score of 0.041 (95% CI: 0.041-0.041), and an expected/observed (E/O) ratio of 1.003 (95% CI: 1.001-1.006), outperforming all three existing risk scores assessed.
+(a) Original existing risk scores.
+(b) Recalibrated existing risk scores.
+Figure 2: Calibration curves.
+Calibration curves for our approach (AP2), QRisk3, SCORE2, and Framingham score, both before and after calibrating the existing risk scores to the UK Biobank cohort.
+Observed risk was calculated using Kaplan-Meier estimators kaplan1958nonparametric .
+QRisk3 and SCORE2 achieved C-indexes of 0.722 (95% CI: 0.721-0.724) and 0.710 (95% CI: 0.709-0.712), respectively, with similar Brier scores (QRisk3: 0.044, 95% CI: 0.043-0.044; SCORE2: 0.042, 95% CI: 0.042-0.42).
+Both approaches overpredicted risk in the UK Biobank population (E/O ratio QRisk3: 1.941, 95% CI: 1.937-1.944; SCORE2: 1.606, 95% CI 1.603-1.608) consistent with previous findings Parsons2023 .
+The Framingham score was the weakest performing risk score, achieving a C-index of 0.705 (95% CI: 0.703-0.707).
+This underperformance likely was a consequence of the Framingham score being developed on a substantially different population from the USA.
+Calibration curves (Fig.
+2) show that AutoPrognosis predictions were well calibrated across all risk deciles.
+In contrast, the existing risk scores overpredicted risk for all individuals in the UK Biobank cohort (Fig.
+2(a)).
+We recalibrated the existing risk scores by scaling the predictions according to the overall incidence in the UK Biobank cohort and the mean predicted risk (Fig.
+2(b)).
+Following recalibration, QRisk3 exhibited good calibration across all risk deciles, while SCORE2 and Framingham score both modestly overpredicted risk in the lower risk deciles and underpredicted risk in the highest risk decile.
+(a) Original existing risk scores.
+(b) Recalibrated existing risk scores.
+Figure 3: Decision curve analysis.
+Our approach (AP2) provides greater net benefit at all thresholds than the existing risk scores (Framingham score, SCORE2, and QRisk3) and the baseline strategies (All and None).
+Figure 4: Feature importance. SHAP values of the variables included in the AutoPrognosis model.
+To assess model performance in a manner that considers the effect on clinical decision making, we additionally performed decision curve analysis Vickers2006 ; Vickers2008 , as recommended in the TRIPOD guidelines Moons2015 .
+Decision curve analysis is a method to evaluate the clinical utility of a model by assessing the net benefit for a range of risk thresholds at which a decision maker would act or intervene.
+The model with the higher net benefit at any given threshold is preferred.
+We compared our model with the existing risk scores, as well as baseline strategies that assume all patients will develop CVD (All) or none will (None).
+Our model achieved a greater net benefit at all decision thresholds compared to the existing risk scores (Fig.
+3).
+Net benefit takes into account both discrimination and calibration.
+While recalibrating the existing risk scores improved their net benefit, the AutoPrognosis model still resulted in the highest net benefit at all thresholds.
+To better understand the rationale for the predictions and drivers of the outperformance of our approach, we assessed feature importance by calculating SHAP values lundberg2017unified .
+SHAP values for the 20 features included in the final version of our model are shown in Fig.
+4.
+Consistent with clinical knowledge, age, sex, and blood pressure are three of the most important features and were included in all three existing risk scores considered in this work.
+However, a number of additional features, including both laboratory and nonlaboratory tests, also significantly contributed to model predictions.
+A number of these features have been shown to be risk factors for CVD, but have not been incorporated into other risk scores.
+For example, Apolipoprotein B (ApoB) is a primary component of several lipoproteins, including low-density lipoprotein (LDL), and transports cholesterol and other lipids in the blood.
+High levels of ApoB are the primary driver of atherosclerosis and there is evidence that ApoB is a more accurate predictor of CVD than total cholesterol or LDL Marston2022 ; Behbodikhah2021 .
+ApoB is deemed one of the most important variables in our approach (Fig.
+4), along with Apolipoprotein A and Lipoprotein A which have also been shown to be risk factors for CVD Erqou2010 .
+Finally, “Overall health rating”, a non-laboratory, self-reported feature, was found to be predictive, mirroring the findings of Alaa et al.
+regarding the benefits of non-laboratory variables for CVD risk prediction Alaa2019 .
 Table 2: Representative questions that a clinician might have relating to a risk score at different stages of a patient encounter, together with whether existing interfaces for risk scores provide this information.
@@
 Stage Representative questions interfaces
-Before Patient
-Which features does the risk score use?
-Encounter
-Why are these features included in the risk score?
-How was the risk score validated?
+Before Which features does the risk score use?
+Patient Why are these features included in the risk score?
+Encounter How was the risk score validated?
 What is the methodology underlying the risk score?
-Before Risk
-When do clinical guidelines recommend risk scoring?
-Scoring
-What is the recommended risk score?
+Before Risk When do clinical guidelines recommend risk scoring?
+Scoring What is the recommended risk score?
 Who is the risk score suitable for?
-During Risk
-What is the risk for this patient?
-Scoring
-What characteristics led to the patient’s risk?
+During Risk What is the risk for this patient?
+Scoring What characteristics led to the patient’s risk?
 What effect would changing this feature have on the risk?
-After Risk Scoring
-What action is recommended by the guidelines based on the risk?
-As a concrete example, let us consider risk scores.
-Currently, clinicians access these tools via fixed user interfaces or application programming interfaces (APIs).
-Numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1), while existing interfaces only serve to calculate risk.
-Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
-By default, LLMs do not possess the ability to access external tools or information.
-Methods to extend the functionality of LLMs beyond text generation are in their infancy but already enable powerful LLM-based systems.
-Below, we describe such a system, illustrated in Fig.
-2.
-Instead of solely relying on an LLM to issue predictions or provide information, we can augment LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM.
-This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface.
-This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and contrasts previous applications of LLMs in medicine.
-1 Illustrative LLM Interfaces
-To illustrate the ideas described in this paper, we built example instantiations of LLM-based interfaces for cardiovascular disease (CVD) and diabetes risk prediction (Fig.
-3).
-To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM.
-Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input.
-This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM.
-Additional implementation details are provided in Appendix A.
-CVD Risk Prediction.
-We first show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig.
-2(a)).
-We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
-Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
-An example interaction between a physician and the LLM-based system is shown in Fig.
-2(a).
-An illustration of the reasoning process by which the LLM uses external tools is provided in Fig.
-S.1).
+After Risk Scoring What action is recommended by the guidelines based on the risk?
 Figure 5: Overview of an LLM-based system that enables clinicians to interface with digital tools using natural language inputs.
@@
 (3) The LLM provides a response to the user incorporating information provided by any tools that were used.
-In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM uses QRisk3 to estimate the patient’s risk, providing this to the user.
-The LLM then summarizes the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines.
-Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges.
+2.2 LLM-based Interfaces
+While LLMs are powerful models for natural language processing, by default, they do not possess the ability to access external tools or information.
+Methods to extend the functionality of LLMs beyond text generation are in their infancy but can already be used to significantly expand the capabilities of LLMs schick2023toolformer ; nakano2021webgpt .
+In this section, we construct LLM-based systems and demonstrate multiple examples of how such systems can provide a novel interface for digital health tools, in particular clinical risk scores.
+The LLM-based systems can incorporate numerous external tools, sources of information, and clinical data (illustrated in Fig.
+5).
+Instead of using an LLM to issue predictions directly, we augmented LLMs and enabled them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM, while unifying multiple digital tools within a single, natural language-based interface.
+To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM.
+Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input.
+This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM.
+Additional implementation details are provided in Section 4.
+Figure 6: Example interactions with LLM-based interfaces.
+(a) Clinicians can ask the LLM questions about the AutoPrognosis risk score, such as what variables are included, the underlying methodology, and the cohort on which the model was derived.
+(b) After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI to improve understanding of the model predictions.
+AutoPrognosis.
+We constructed an LLM-based interface for the machine learning model developed in this work (Fig.
+6).
+In our first example, we demonstrate the role such an interface could have upon model deployment (Fig.
+6(a).
+A clinician might have many questions about a risk score before using the predictions to inform recommendations or interventions (Table 2).
+Here, the LLM provided specific information about the variables used by the risk score, the underlying methodology, how the model was constructed, and the cohort of individuals used to train the model.
+Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 .
+A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
+To enable this, in addition to allowing the LLM to access the model, we provided the LLM with the ability to apply methods from explainable AI (XAI) Imrie2023 to the underlying model, allowing users to investigate the rationale for predictions.
+Figure 7: Example interaction with an LLM-based interface for QRisk3.
+The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines.
+As shown in Fig.
+6(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model.
+In this example, we used Shapley additive explanations (SHAP) lundberg2017unified to explain the prediction.
+For this individual, their estimated 4.2% risk was primarily caused by their age and family history of CVD, mitigated by being a woman, not smoking, and low levels of lipoprotein (a).
+An additional interaction can be found in Fig.
+S.3.
+Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM.
+For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified .
+The LLM could explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified.
+This demonstrates the utility of LLMs beyond simply using existing tools and the benefit of their underlying knowledge.
+QRisk3.
+We also show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig.
+7).
+We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
+Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
+An example interaction between a physician and the LLM-based system is shown in Fig.
+7.
+An illustration of the reasoning process by which the LLM uses external tools is provided in Fig.
+S.1.
+In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM used QRisk3 to estimate the patient’s risk, providing this to the user.
+The LLM then summarized the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines.
+Finally, the LLM used QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges.
 This allows both the clinician and patient to understand the potential impact of changes to modifiable variables on the patient’s risk.
 While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 .
-In the following example, we will explore a more dynamic output extending beyond model use and information retrieval to new functionality.
-(a) Cardiovascular disease risk prediction
-(b) Diabetes risk prediction
-Figure 3: Example interactions with LLM-based interfaces.
-(a) CVD: The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines.
-(b) Diabetes: After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
-Diabetes Risk Prediction.
-We now demonstrate an application for diabetes risk prediction using a machine learning model (Fig.
-2(b)).
-We trained a prognostic model using AutoPrognosis Imrie2023AutoPrognosis , an open-source automated machine learning framework, on a cohort of individuals from UK Biobank Sudlow2015 to estimate a patient’s 10-year risk of developing diabetes.
-Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 .
-A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
-To enable this, in addition to allowing the LLM access to the model, we provided the LLM with the ability to apply explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions.
-As shown in Fig.
-2(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model.
-In this example, we demonstrate the use of Shapley additive explanations (SHAP) lundberg2017unified .
-Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM.
-For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified .
-The LLM is able to explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified.
-This demonstrates the utility of LLMs beyond simply using existing tools and illustrates the benefit of their underlying knowledge.
 3 Discussion
 Large Language Models hold substantial promise for the medical domain, particularly in augmenting digital workflows and improving efficiency in healthcare delivery.
-The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as the potential to hallucinate.
+The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as hallucinations.
+Doing so offers a potential transformation for how clinicians interact with digital tools and could help overcome the challenges of deploying clinical AI models.
 We have demonstrated how LLMs can provide a unique interface between healthcare professionals and clinical predictive models, such as risk scores.
-Our approach is scalable and requires no additional training, although approaches that improve with use could be yet more powerful.
-Additionally, we aim to mitigate the problem of LLM hallucination and ensure that actionable advice is anchored in approved clinical resources.
+In particular, we developed a novel risk score for incidence of CVD using automated machine learning and developed LLM-based interfaces for our model and QRisk3 Hippisley-Cox2017CVD , the current recommended risk score in the UK for CVD.
+Our approach is scalable and does not require any additional training of the language model, although approaches that improve with use could be yet more powerful.
+Additionally, we aim to mitigate the problem of hallucination by ensuring that actionable advice is anchored in approved clinical resources, contrasting several previous applications of LLMs in medicine that focused exclusively on the knowledge and information learned by LLMs.
+Numerous considerations exist before, during, and after assessing an individual’s prognosis (Table 2).
+Currently, clinicians must access these tools via fixed user interfaces or application programming interfaces (APIs), with existing interfaces typically only calculating risk.
+Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
+In this paper, we have focused on clinicians interacting with digital tools.
+However, there are numerous stakeholders in healthcare in addition to clinicians, such as patients, regulators, and administrators, each with different goals and requirements Imrie2023 .
+Our framework and approach could be applied to improve digital health interfaces for these alternate stakeholders.
+While this could have additional challenges, there are potentially even more substantial benefits for such individuals, given the differences in requirements, knowledge, and familiarity with digital health technology, among other factors.
 While LLMs have general capabilities, they can lack domain-specific knowledge.
 This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 .
-While we showed using such specialist LLMs is not required, they are easy to incorporate due to the modularity of the approach and doing so could further enhance the functionality of LLM-based interfaces.
-As AI in medicine continues to advance, further research into LLMs and their potential applications is not only beneficial but vital.
-LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes.
+While we showed using such specialist LLMs is not required, they could be readily incorporated due to the modularity of our approach.
+This could further enhance the functionality of LLM-based interfaces.
+As AI in medicine continues to advance, further research into LLMs and their potential applications in healthcare could provide significant benefits.
+For example, LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes.
 Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment.
@@
 Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
+4 Methods
+4.1 Data source and study population
+We developed and validated our prognostic model using data from UK Biobank Sudlow2015 , a large prospective cohort of individuals from the UK.
+Participants in UK Biobank were enrolled between 2006 and 2010 and aged between 40 and 70 at the time of recruitment.
+We extracted a cohort of individuals with no known history of CVD at baseline.
+We excluded individuals who had been diagnosed with chronic kidney disease or type 1 diabetes and individuals who were being treated with statins.
+This resulted in a cohort of 407,605 individuals (Fig.
+S.2).
+4.2 Clinical Predictors and Outcomes
+We began by considering a number of predictors previously validated as CVD risk factors.
+A subset of features were selected based on their predictive power and feature importance.
+We included 20 features in our model, namely sex, age, atrial fibrillation, steroid treatment, type 2 diabetes, hypertension treatment, family history of CVD, systolic blood pressure, smoking status (never-smoker, ex-smoker, light smoker, moderate smoker, heavy smoker, alkaline phosphatase, apolipoprotein a, apolipoprotein b, cystatin c, C-reactive protein, HbA1c, IGF-1, lipoprotein a, triglycerides, urea, and overall health rating, which was self-reported by participants as excellent, good, fair or poor.
+The primary outcome was incidence of CVD within a 10-year horizon.
+This was defined using the following ICD-10 codes: myocardial infarction (I21, I22), stroke (I63, I64), angina pectoris (I20), or transient cerebral ischaemic attacks (G45).
+There were 17,600 CVD events within a 10-year horizon in the UK Biobank cohort.
+4.3 Model Derivation
+We trained a prognostic model using AutoPrognosis 2.0, an open-source automated machine learning software package Imrie2023AutoPrognosis .
+AutoPrognosis has been validated in several applications in medicine, for example to determine eligibility for lung cancer screening callender2023assessing .
+AutoPrognosis was used to optimize pipelines consisting of a variable preprocessing step followed by model selection and training.
+The optimized pipelines were subsequently combined in an ensemble using a weighted combination.
+Individual pipelines and the final ensemble were trained to maximize discrimination, measured using the area under the receiver operating curve.
+Predictions from the final ensemble were calibrated in a piecewise linear manner using the observed risk calculated using a Kaplan-Meier estimator kaplan1958nonparametric .
+4.4 Statistical Analysis
+Model discrimination was assessed using the concordance index (C-index) Uno2011 and calibration via calibration curves and the ratio between the expected and observed number of events.
+We also assessed Brier scores brier1950verification .
+Calibration curves were constructed by first dividing the population into 10 risk deciles based on each individual’s predicted 10-year risk and then comparing the mean estimated risk to the observed risk in each decile.
+Kaplan–Meier estimators kaplan1958nonparametric were used to calculate observed risk.
+Additionally, we conducted decision curve analysis, assessing the net benefit for a range of risk thresholds.
+In addition to the existing risk scores, when assessing net benefit, we also compared our model with baseline strategies that assume all patients develop CVD (All) or none do (None).
+Results are based on 5-fold cross-validation.
+We used multiple imputation with chained equations (MICE) vanBuuren2011 to impute missing values.
+We generated five imputed datasets and combined results using Rubin’s rules rubin2018multiple .
+4.5 Model Comparisons
+We compared our model with three existing risk scores, namely Framingham score DAgostino2008 , SCORE2 SCORE2 , and QRisk3 Hippisley-Cox2017CVD .
+All three models assess the likelihood of developing CVD within a 10 year horizon.
+The Framingham score DAgostino2008 was recommended by the 2010 American College of Cardiology/American Heart Association (ACC/AHA) guidelines Greenland2010 and is based on eight risk factors: sex, age, systolic blood pressure, treatment for hypertension, smoking status, history of diabetes, total cholesterol, and high-density lipoprotein (HDL) cholesterol.
+SCORE2 SCORE2 was developed using data from 45 cohorts in 13 countries across Europe, including the United Kingdom, and has been included in the European Society of Cardiology guidelines.
+SCORE2 employs seven risk factors: sex, age, systolic blood pressure, smoking status, history of diabetes, total cholesterol, and high-density lipoprotein (HDL) cholesterol.
+QRisk3 Hippisley-Cox2017CVD was developed on data from 981 practices in England comprising almost 8 million patients and is currently the recommended risk score for CVD in England NICE_CVD .
+QRisk3 uses 22 variables, including demographics (such as sex, age), lifestyle factors (e.g.
+smoking), medical and family history (e.g.
+history of diabetes, family history of CVD), medical examination (e.g.
+blood pressure), and blood tests (total cholesterol/HDL cholesterol ratio).
+4.6 LLM-based interface
+Here, we describe the implementation details for the LLM-based systems demonstrated above.
+4.6.1 LLM framework
+By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history.
+To construct interfaces for digital tools using LLMs, we instead viewed the LLM as an agent that can interact with an environment to solve tasks.
+Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history.
+To enable the agent to both reason and use external tools, we used the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
+Since we will provide the agents tasks in the form of natural language, and actions in the language space $L$ are (essentially) infinite, we chose to benefit from strong language priors and use a pretrained LLM.
+We implemented our LLM-based interfaces using GPT-4 openai2023gpt4 , with access through the OpenAI API.
+Interactions with external tools were implemented using LangChain LangChain .
+Frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively.
+In contrast, following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM.
+This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples.
+This also enabled us to use multiple tools.
+We built user interfaces using StreamLit streamlit .
+4.6.2 External Tools
+This section describes the external tools and sources of information made available to the LLM in our examples.
+LLM-based interface for AutoPrognosis.
+To demonstrate an LLM-based interface, we equipped the LLM with the CVD risk prediction model developed in this work on the UK Biobank cohort using AutoPrognosis 2.0 Imrie2023AutoPrognosis .
+Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial .
+A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
+To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient.
+In particular, we used SHAP lundberg2017unified to interpret model predictions.
+We additionally provided the LLM with information about the variables used by the risk score, the underlying methodology and how the model was constructed, and details regarding the cohort used to develop the model.
+Example interactions are shown in Fig.
+6 and Fig.
+S.3.
+LLM-based interface for QRisk3.
+As a second example, we showed how LLMs can incorporate existing tools and information for CVD risk prediction.
+We provided the LLM access to QRisk3 Hippisley-Cox2017CVD and enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk.
+Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
+An example interaction is shown in Fig.
+7.
 5 Code availability
 Code for the LLM-based interfaces can be accessed at https://github.com/pauliusrauba/LLMs_interface.
+AutoPrognosis is an open-source package available on GitHub at https://github.com/vanderschaarlab/AutoPrognosis and on PyPI at https://pypi.org/project/autoprognosis/.
 6 Data availability
-The risk predictor for diabetes was trained on a cohort from the UK Biobank resource.
+This research has been conducted using the UK Biobank resource under application number 105160.
 Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register).
@@
 wrote the original draft of the manuscript, and all authors contributed to editing and revising it.
-    A.1 Large Language Models
-    By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history.
-To construct interfaces for digital tools using LLMs, we instead view the LLM as an agent that can interact with an environment to solve tasks.
-Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history.
-To enable the agent to both reason and use external tools, we use the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
-    Since we will provide the agents tasks in the form of natural language, and actions in the language space $L$ are (essentially) infinite, we chose to benefit from strong language priors and use a pretrained LLM.
-We implemented our LLM-based interfaces using GPT-4 openai2023gpt4 , with access through the OpenAI API.
-Interactions with external tools were implemented using LangChain LangChain .
-    While frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively.
-Following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM.
-This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples.
-This also enabled us to use multiple tools.
-We built user interfaces using StreamLit streamlit .
-    A.2 External Tools
-    This section describes the external tools and sources of information made available to the LLM in our two examples. In the subsequent section, we provide the system and tool prompts.
-    CVD Risk Prediction.
-In our first example, we show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig.
-3a).
-We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years.
-QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD .
-We enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk.
-Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
-    Diabetes Risk Prediction.
-To demonstrate an application for diabetes risk prediction using a machine learning model (Fig.
-3b), we equipped the LLM with the following tools.
-For the diabetes risk score, we trained a predictive model using AutoPrognosis 2.0 Imrie2023AutoPrognosis on a cohort of individuals from UK Biobank Sudlow2015 with no known history of diabetes at baseline.
-The primary outcome was diagnosis of diabetes within a 10-year horizon.
-We selected 20 features based on their predictive value Imrie2023AutoPrognosis , namely HbA1c, blood glucose, sex, age, waist size, hip size, weight, BMI, waist-height ratio, waist-hip ratio, gamma-glutamyl transferase, cystatin-c, C-reactive protein, atrial fibrillation, HDL, alanine transaminase, triglycerides, uric acid, sex hormone-binding globulin, use of anti-hypertensives.
-    Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial .
-A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking .
-To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient.
-In particular, we used SHAP lundberg2017unified to interpret model predictions.
-    Figure S.1: Illustration of the process by which the LLM uses external tools.
-The LLM is provided with the history of the interaction, including the current request.
-Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input (“Thought”).
-The LLM then uses the external tool (“Action”) and receives the output (“Observation”).
-Finally, the LLM decides to answer the question (“Thought”), providing a response to the user.